OpenStudy (anonymous):

If line mn is the diameter of a circle with M(-3, -4) and N(-7, -14), find the center of the circle

OpenStudy (campbell_st):

the centre of the circle will be the midpoint of the diameter. so find the midpoint of (-3, -4) and (-7, -14)

OpenStudy (anonymous):

Midpoint: \[(\frac{ x _{2} +x _{1}}{ 2 }, \frac{ y _{2} +y _{1}}{ 2 })\] = ((-7-3)/2,((-14-4)/2) = (-5,-9)

OpenStudy (anonymous):

Center is (-5,-9)

OpenStudy (campbell_st):

why give the answer... just give the formula for the midpoint. Let the asker work through the information